OpenStudy (anonymous):

Help! How many moles of a gas sample are in 2.5 L container at 373 K and 203 kPa?

OpenStudy (adamconner):

do you know the gas law?

OpenStudy (adamconner):

the ideal gas law, which you will need to memorize is : PV=nRT pressure in atm volume in liters n is mole R constant is 0.0821 L*atm/mol*K T in kelvin

OpenStudy (adamconner):

you have pressure in kPa, change it to atm and you will get 2.003 this way everything else is ready, just plug and chug

OpenStudy (adamconner):

PV=nRT (2.003)(2.5)=n(.0821)(373) 5.0075=n(30.6233) 5.0075/30.6233=n n=.1635 mol